Write in slope-intercept form (if possible) and standard form an equation of the line described
slope = 0 and contains (-2, -5)

Answer by mananth(16946)

y= mx +b slope intercept form
m=0, (-2,-5)
y=0*(-2)-5
so y=-5 is the equation

--
y+5 =0 standard form